# Integrated Sensing and Communications in Downlink FDD MIMO without CSI Feedback

### v1

@@ Abstract
In this paper, we propose a precoding framework for frequency division duplex (FDD) integrated sensing and communication (ISAC) systems with multiple-input multiple-output (MIMO). Specifically, we aim to maximize ergodic sum spectral efficiency (SE) while satisfying a sensing beam pattern constraint defined by the mean squared error (MSE). Our method reconstructs downlink (DL) channel state information (CSI) from uplink (UL) training signals using partial reciprocity, eliminating the need for CSI feedback. To mitigate interference caused by imperfect DL CSI reconstruction and sensing operations, we adopt rate-splitting multiple access (RSMA). We observe that the error covariance matrix of the reconstructed channel effectively compensates for CSI imperfections, affecting both communication and sensing performance. To obtain this, we devise an observed Fisher information-based estimation technique. We then optimize the precoder by solving the Karush-Kuhn-Tucker (KKT) conditions, jointly updating the precoding vector and Lagrange multipliers, and solving the nonlinear eigenvalue problem with eigenvector dependency to maximize SE. The numerical results show that the proposed design achieves precise beam pattern control, maximizes SE, and significantly improves the sensing-communication trade-off compared to the state-of-the-art methods in FDD ISAC scenarios.
@@
Integrated sensing and communications (ISAC), rate-splitting multiple access (RSMA), error covariance estimation, generalized power iteration (GPI).
@@ I Introduction
The convergence between wireless communication and sensing technologies is driving the development of innovative applications in next-generation networks, such as autonomous driving, smart cities, and environmental monitoring systems [1, 2, 3]. In this regard, the concept of integrated sensing and communication (ISAC) has been recognized as a vital approach, enabling simultaneous sensing and communication within a unified system architecture [1]. This method not only improves spectral efficiency (SE), but also significantly reduces hardware costs and energy consumption by sharing resources between sensing and communication [4, 5].
Integrating ISAC into current cellular systems requires a critical decision on the duplex mode, choosing between frequency division duplex (FDD) and time division duplex (TDD) [6, 1, 5]. Recent studies show that while TDD ISAC benefits from simplified channel estimation thanks to channel reciprocity, it may not be suitable for low-latency services due to the inherent delay caused by UE and BS having to wait for their respective transmission slots [7, 5, 6]. Additionally, TDD ISAC requires an extended guard interval to receive and process echo signals for target detection before uplink (UL) transmission, further increasing latency. This poses challenges for applications sensitive to delays, such as autonomous driving [7, 5]. As a remedy, FDD allows simultaneous UL and DL transmissions, eliminating the need for subframe switching and guard intervals. This underscores FDD’s importance for low-latency ISAC applications. The relationship between ISAC echo signals and UL communication signals for each duplex mode is shown in Fig. 1. Unfortunately, however, FDD still faces challenges due to the lack of channel reciprocity, which leads to substantial overhead in acquiring DL CSI for communication, which scales with the number of antennas and users [8, 9]. To address this, we propose a novel precoding framework that reconstructs DL CSI from UL training and enhances both communication and sensing capabilities.
@@ I-A Related Work
Early studies on the coexistence of communication and radar systems, which operated on separate hardware with shared spectrum [10], paved the way for the development of ISAC based on a unified infrastructure. In ISAC systems, maximizing efficiency requires the simultaneous support of multiple communication users and detection targets. This has driven research into MIMO-ISAC.
@@
In the literature on MIMO-ISAC, the design of the transmit precoder has been a key topic. For example, [11] developed a branch-and-bound framework to achieve the global optimal solution that minimizes the mean square error (MSE) of the beam pattern under a constant modulus constraint. However, this approach does not assume designated radar symbols, making it sub-optimal compared to considering respective symbols for communication and sensing. To address this, a joint beamforming design for dedicated communication and sensing symbols was proposed in [3]. The authors in [12] introduced a generalized Pareto optimization framework, applying a bisection search algorithm to meet the lowest acceptable performance levels for dual functionalities. More recently, [4] proposed a rate-splitting multiple access (RSMA)-based precoder optimization technique to mitigate interference caused by added sensing functionality. However, the studies in [11, 3, 12, 4] did not account for the imperfections in CSI at the transmitter (CSIT), which are inevitable in practical wireless communication systems [13, 14].
Figure 1: While TDD ISAC, which uses the same bandwidth for DL and UL, requires a sufficient guard interval to separate the echo signal from the UL communication, FDD ISAC does not require a guard interval due to its separation in the frequency domain.
@@
In light of this, there have also been research endeavors that take into account CSIT imperfection. Specifically, in [15], channel estimation errors were addressed for MIMO-ISAC, whose key idea is minimizing the worst-cast multi-user interference. In [16, 17], robust beamforming designs for ISAC were proposed using techniques related to physical layer security or reconfigurable intelligent surface (RIS). In addition, robust ISAC beamforming methods for coping with CSIT imperfection were also investigated in [18, 16, 17, 2]. In this work, two types of CSIT error modeling were adopted:
@@
Gaussian error with known covariance : The CSIT error is modeled as a Gaussian with known covariance [18, 16, 2]. This corresponds to FDD MIMO with scalar quantization.
@@
Bounded error: The magnitude of the CSIT error is assumed to be bounded within certain ranges [17, 16, 15]. This corresponds to FDD MIMO with vector quantization.
@@
In summary, the existing error modeling is established under the assumption of channel feedback in FDD MIMO systems. However, we recall that the amount of feedback often negates SE gains, as it scales with the number of antennas and users [8, 14], which fundamentally limits the potential gains of MIMO-ISAC; thus developing a low-complexity CSI acquisition approach is essential. To address this issue, in [19, 20, 21, 22], the concept of eliminating CSI feedback was explored in FDD MIMO to simplify the CSI acquisition process, known as zero-feedback strategies. A key idea of the zero-feedback strategy is to leverage partial channel reciprocity to efficiently reconstruct the DL channel from UL reference signals, by which we can omit the direct channel feedback even in FDD. However, with these DL CSIT reconstruction approaches, the CSIT error cannot be avoided due to the discrepancies between the UL and DL environments [20]. Unfortunately, it is not feasible to employ the previous CSIT error treatment approaches to handle the CSIT error caused by zero-feedback strategies [20, 22]. That is to say, it is reasonable to assume that the corresponding CSIT error is not bounded as in [17, 16, 15] since CSI quantization is not used in zero-feedback strategies. Further, its error covariance cannot be obtained in a straightforward way because highly nonlinear algorithms are typically used to reconstruct the DL CSIT from UL reference signals such as [23]. To resolve this, there have been attempts to approximate the CSIT error covariance using the Cramér-Rao lower bound matrix (CRLB) [24, 22]. Specifically, in [24], they computed the CRLB matrix in an instantaneous manner during UL training. Incorporating this into the precoder design, it was found that significant SE gains could be achieved even without CSI feedback in FDD MIMO. Consequently, it is promising to explore the zero-feedback strategy for MIMO-ISAC.
@@ I-B Contributions
In this paper, we present a novel FDD MIMO-ISAC transmission framework that does not necessitate CSI feedback. Unlike a conventional approach, we reconstruct the DL CSI solely based on UL sounding reference signals by following [21, 24]. In this setup, acquiring perfect DL CSIT is infeasible, making efficient mitigation of interference arising from imperfect DL CSIT crucial to achieve robust SE and ensure sensing performance. In this perspective, we summarize our main contributions as follows.
  • We investigate a DL channel reconstruction method in which channel parameters are estimated solely from UL training. The main problem is then formulated to maximize the sum SE under imperfect CSIT while meeting a beam pattern matching constraint based on a target MSE. To address CSIT inaccuracies caused by zero feedback, we consider the RSMA technique. The communication rate for the common message is defined using a minimum function to ensure that it is decodable by all users. For the sensing constraint, we represent the MSE between the desired and actual beam pattern in a quadratic form. In solving the formulated problem, we find that accurate CSI error covariance matrix estimation is crucial to designing precoders that effectively cope with CSIT inaccuracies.
@@ II-A Communication Channel Model
where $\beta_{k,\ell}$ is assumed to be a random variable following $\mathcal{CN}(0,\sigma^{2}_{\mathrm{path},k}),\forall\ell$, independent of the UL channel gain $\alpha_{k,\ell}^{\mathrm{ul}}$. In the special case where $\eta_{k,\ell}=1,\forall(k,\ell)$, our model reduces to a reciprocal channel gain case ($\alpha_{k,\ell}^{\mathrm{dl}}=\alpha_{k,\ell}^{\mathrm{ul}}$) considered in the previous work [19, 26], and therefore we consider a more generalized scenario.
@@ II-B RSMA-assisted MIMO-ISAC Signal Model
In this study, we anticipate that the unified hardware structure for ISAC may lead to a lack of available spatial resources. To effectively address this, we consider RSMA, a multiple access technique that enhances robustness when spatial resources are shared for multiple objectives [28]. Specifically, RSMA mitigates interference between sensing and communication, as well as among communication users [4, 29, 25]. In RSMA, each user’s message $M_{k}$ is split into a common part $M_{\mathrm{c},k}$ and a private part $M_{\mathrm{p},k}$. The common parts are encoded into a single common stream $s_{\mathrm{c}}$, while private parts are encoded individually as $s_{k}$, $k\in[K]$. At the receiver, each user $k$ first decodes the common stream, removes it using successive interference cancellation (SIC), and then decodes the private stream $s_{k}$ while treating residual interference as noise [28].
@@
where $\mathbf{p}_{\mathrm{c}}$ denotes the precoding vector for the common message symbol $s_{\mathrm{c}}$, $\mathbf{P}_{\mathrm{p}}$ is a precoding matrix for private message vector $\mathbf{s}_{\mathrm{p}}=[s_{\mathrm{p},1},s_{\mathrm{p},2},\ldots,s_{\mathrm{p},K}]^{\sf T}$, which is defined as $\mathbf{P}_{\mathrm{p}}=[\mathbf{p}_{\mathrm{p},1},\mathbf{p}_{\mathrm{p},2},\ldots,\mathbf{p}_{\mathrm{p},K}]\in\mathbb{C}^{N\times K},$ where $\mathbf{p}_{\mathrm{p},k}$ is the precoder for the private message of user $k$. Likewise, $\mathbf{P}_{\mathrm{r}}$ represents the precoding matrix for the vector of radar sequence symbols $\mathbf{s}_{\mathrm{r}}=[s_{\mathrm{r},1},s_{\mathrm{r},2},\ldots,s_{\mathrm{r},M}]^{\sf T}$, that is, $\mathbf{P}_{\mathrm{r}}=[\mathbf{p}_{\mathrm{r},1},\mathbf{p}_{\mathrm{r},2},\ldots,\mathbf{p}_{\mathrm{r},M}]\in\mathbb{C}^{N\times M}$, where $\mathbf{p}_{\mathrm{p},m}$ is the radar beamformer for the $m$-th radar symbol. For notational simplicity, we integrate the common and private part with a single expression, which is given by $\mathbf{P}_{\mathrm{c}}=[\mathbf{p}_{\mathrm{c}},\mathbf{P}_{\mathrm{p}}]$ and $\mathbf{s}_{\mathrm{c}}=[s_{\mathrm{c}},\mathbf{s}_{\mathrm{p}}^{\sf T}]^{\sf T}$. Lastly, we define the entire precoding vector and the symbol vector as $\mathbf{P}=[\mathbf{P}_{\mathrm{c}},\mathbf{P}_{\mathrm{r}}]$ and $\mathbf{s}=[\mathbf{s}_{\mathrm{c}}^{\sf T},\mathbf{s}_{\mathrm{r}}^{\sf T}]^{\sf T}$, respectively. The communication symbols $\mathbf{s}_{\mathrm{c}}$ and the radar sequences $\mathbf{s}_{\mathrm{r}}$ are statistically independent of each other [2], that is, $\mathbb{E}[\mathbf{s}_{\mathrm{c}}\mathbf{s}_{\mathrm{r}}^{\sf H}]=\mathbf{0}_{(K+1)\times M}$. In addition, the power constraint with transmit power $P$ is represented by $\mathbb{E}[{\mathbf{s}_{\mathrm{c}}\mathbf{s}_{\mathrm{c}}^{\sf H}}]=P\cdot\mathbf{I}_{K+1}$ and $\mathbb{E}[{\mathbf{s}_{\mathrm{r}}\mathbf{s}_{\mathrm{r}}^{\sf H}}]=P\cdot\mathbf{I}_{M}$, respectively. In this regard, we presume that $\|\mathbf{P}_{\mathrm{c}}\|_{\sf F}+\|\mathbf{P}_{\mathrm{r}}\|_{\sf F}\leq 1$, without loss of generality.

### v3

@@ Abstract
In this paper, we propose a precoding framework for frequency division duplex (FDD) integrated sensing and communication (ISAC) systems with multiple-input multiple-output (MIMO). Specifically, we aim to maximize ergodic sum spectral efficiency (SE) while satisfying a sensing beam pattern constraint defined by the mean squared error (MSE). Our method reconstructs downlink (DL) channel state information (CSI) from uplink (UL) training signals using partial reciprocity, eliminating the need for CSI feedback. To obtain the error covariance matrix of the reconstructed DL CSI, we devise an observed Fisher information-based estimation technique. Leveraging this, to mitigate interference caused by imperfect DL CSI reconstruction and sensing operations, we propose a rate-splitting multiple access (RSMA) aided precoder optimization method. This method jointly updates the precoding vector and Lagrange multipliers by solving the nonlinear eigenvalue problem with eigenvector dependency to maximize SE. The numerical results show that the proposed design achieves precise beam pattern control, maximizes SE, and significantly improves the sensing-communication trade-off compared to the state-of-the-art methods in FDD ISAC scenarios.
@@
Integrated sensing and communications (ISAC), rate-splitting multiple access (RSMA), error covariance estimation, Non-linear eigenvalue problem.
@@ I Introduction
The convergence of wireless communication and sensing technologies is accelerating the development of intelligent systems such as autonomous driving, smart cities, and environmental monitoring [1, 2, 3]. A key enabler of these applications is integrated sensing and communication (ISAC), which allows both functionalities to share infrastructure and spectrum within a unified architecture. ISAC not only improves spectral efficiency but also reduces hardware complexity and energy consumption [1, 4, 5].
When deploying ISAC in cellular systems, one of the most critical design choices is the duplexing mode: time division duplex (TDD) or frequency division duplex (FDD) [6, 1, 5]. TDD systems can benefit from channel reciprocity, which simplifies downlink (DL) channel estimation. However, enabling monostatic sensing in TDD typically requires dedicating separate time slots for sensing to avoid collisions between reflected sensing signals and uplink (UL) communication, which is incompatible with current communication protocols [5]. This leads to implementation challenges and may introduce additional latency [6]. In contrast, FDD systems allocate separate frequency bands for UL and DL transmissions, enabling simultaneous communication and sensing without interference, as shown in Fig. 1. This allows ISAC functionality without modifications to existing communication standards [5, 1]. Moreover, FDD is already widely deployed in cellular networks, making it a practical option for near-term ISAC deployment.
To meet the increasing demands of ISAC—such as supporting multiple users, enabling high-resolution sensing, and managing interference—multiple-input multiple-output (MIMO) technology is essential. MIMO provides rich spatial degrees of freedom, allowing advanced beamforming and spatial multiplexing to support both communication and sensing tasks effectively [1, 5]. Within this context, transmit precoding plays a crucial role in managing multi-user interference (MUI) and jointly optimizing the performance of both functionalities [2, 4]. Nonetheless, a fundamental challenge remains: FDD systems lack channel reciprocity, so DL channel state information (CSI) must be obtained through feedback, resulting in non-trivial overhead and potentially degraded performance in dynamic environments [7, 8].
In this paper, we propose a novel precoding framework for FDD-based massive MIMO-ISAC systems that eliminates the need for CSI feedback by reconstructing DL CSI from UL training signals. Our approach integrates rate-splitting multiple access (RSMA) with error covariance matrix (ECM) estimation to mitigate MUI and also enhance the sensing performance. The proposed framework is particularly suitable for delay-sensitive applications such as smart cities and autonomous driving, where reducing latency and signaling overhead is crucial [6].
@@ I-A Related work
In the literature on MIMO-ISAC, the design of the transmit precoder has been a key topic. For example, [9] developed a branch-and-bound framework to achieve the global optimal solution that minimizes the mean square error (MSE) of the beam pattern under a constant modulus constraint. However, this approach does not assume designated radar symbols, making it sub-optimal compared to considering respective symbols for communication and sensing. To address this, a joint beamforming design for dedicated communication and sensing symbols was proposed in [3]. The authors in [10] introduced a generalized Pareto optimization framework, applying a bisection search algorithm to meet the lowest acceptable performance levels for dual functionalities. A recent study [11] proposed a hybrid radar fusion (HRF) scheme for FDD systems, combining monostatic and bistatic sensing to enhance angle-of-arrival (AoA) estimation. [4] proposed a RSMA-based precoder optimization technique to mitigate interference caused by added sensing functionality. However, the studies in [9, 3, 10, 4] did not account for the imperfections in CSI, which are inevitable in practical wireless communication systems [12, 13].
Figure 1: TDD ISAC may require modified time slot allocation for signal separation [14]; otherwise, interference between reflected sensing and UL communication signals may occur. In contrast, FDD ISAC leverages separate DL and UL frequency bands to achieve signal separation using established protocols.
@@
In light of this, there have also been research endeavors that take into account CSI imperfection. Specifically, in [15], channel estimation errors were addressed for MIMO-ISAC, whose key idea is minimizing the worst-case multi-user interference. In [16, 17], robust beamforming designs for ISAC were proposed using techniques related to physical layer security or reconfigurable intelligent surface (RIS). In addition, robust ISAC beamforming methods for coping with CSI imperfection were also investigated in [18, 16, 17, 2]. In these works, two types of CSI error modeling were adopted:
@@
Gaussian error with known covariance : The CSI error is modeled as a Gaussian with known covariance [18, 16, 2]. This corresponds to FDD MIMO with scalar quantization.
@@
Bounded error: The magnitude of the CSI error is assumed to be bounded within certain ranges [17, 16, 15]. This corresponds to FDD MIMO with vector quantization.
@@
In summary, the existing error modeling is established under the assumption of channel feedback in FDD MIMO systems. However, we recall that the amount of feedback often negates SE gains, as it scales with the number of antennas and users [7, 13], which fundamentally limits the potential gains of MIMO-ISAC; thus developing a low-complexity CSI acquisition approach is essential. To address this issue, in [19, 20, 21, 22], the concept of eliminating CSI feedback was explored in FDD MIMO to simplify the CSI acquisition process, known as zero-feedback strategies. A key idea of the zero-feedback strategy is to leverage partial channel reciprocity to efficiently reconstruct the DL channel from UL reference signals, by which we can omit the direct channel feedback even in FDD. However, with these DL CSI reconstruction approaches, the CSI error cannot be avoided due to the discrepancies between the UL and DL environments [20]. Unfortunately, it is not feasible to employ the previous CSI error treatment approaches to handle the CSIT error caused by zero-feedback strategies [20, 22]. That is to say, unlike the approaches in [17, 16, 15], where CSI quantization ensures a bounded CSI error, the CSIT error in zero-feedback strategies cannot be assumed to be bounded. Further, its error covariance cannot be obtained in a straightforward way because highly nonlinear algorithms are typically used to reconstruct the DL CSI from UL reference signals such as [23]. To resolve this, there have been attempts to approximate the CSI error covariance using the Cramér-Rao lower bound (CRLB) matrix [24, 22]. Specifically, in [24], they computed the CRLB matrix in an instantaneous manner during UL training. Incorporating this into RSMA precoder design, significant SE gains were achieved in FDD MIMO systems without requiring CSI feedback. These findings highlight the potential of a zero-feedback strategy for MIMO-ISAC systems.
@@ I-B Contributions
In this paper, we present a novel FDD MIMO-ISAC transmission framework that does not necessitate CSI feedback. Unlike a conventional approach, we reconstruct the DL CSI solely based on UL sounding reference signals by following [21, 24]. In this setup, acquiring perfect DL CSI is infeasible, making efficient mitigation of interference arising from imperfect DL CSI crucial to achieve robust SE and ensure sensing performance. In this perspective, we summarize our main contributions as follows.
  • We investigate a DL channel reconstruction method in which channel parameters are estimated solely from UL training. The main problem is then formulated to maximize the sum SE under imperfect DL CSI while meeting a beam pattern matching constraint based on a target MSE. To address CSI inaccuracies caused by zero feedback, we consider the RSMA technique. The communication rate for the common message is defined using a minimum function to ensure that it is decodable by all users. For the sensing constraint, we represent the MSE between the desired and actual beam pattern in a quadratic form. In solving the formulated problem, we find that accurate CSI error covariance matrix estimation is crucial to designing precoders that effectively cope with CSI inaccuracies.
@@ II-A Communication channel model
where $\beta_{k,\ell}$ is assumed to be a random variable following $\mathcal{CN}(0,\sigma^{2}_{\mathrm{path},k}),\forall\ell$, independent of the UL channel gain $\alpha_{k,\ell}^{\mathrm{ul}}$. We consider a more generalized channel model compared to previous works [19, 26], which assumed perfectly reciprocal channel gains and fixed angular parameters. In contrast, our model allows non-reciprocal channel gains ($\eta_{k,\ell}\neq 1$) and accommodates variations in angular characteristics across UL and DL, thus capturing practical deviations in FDD systems more accurately.
@@ II-B RSMA-assisted MIMO-ISAC signal model
In this study, we notice that the unified ISAC architecture may restrict the spatial channel resources between the user and BS. To effectively address this, we consider RSMA, a multiple access technique that enhances robustness when spatial resources are shared for multiple objectives [28]. Specifically, RSMA mitigates interference between sensing and communication, as well as among communication users [4, 29, 25]. In RSMA, each user’s message $M_{k}$ is split into a common part $M_{\mathrm{c},k}$ and a private part $M_{\mathrm{p},k}$. The common parts are encoded into a single common stream $s_{\mathrm{c}}$, while private parts are encoded individually as $s_{k}$, $k\in[K]$. At the receiver, each user $k$ first decodes the common stream, removes it using successive interference cancellation (SIC), and then decodes the private stream $s_{k}$ while treating residual interference as noise [28].
@@
where $\mathbf{p}_{\mathrm{c}}$ denotes the precoding vector for the common message symbol $s_{\mathrm{c}}$, $\mathbf{P}_{\mathrm{p}}$ is a precoding matrix for private message vector $\mathbf{s}_{\mathrm{p}}=[s_{\mathrm{p},1},s_{\mathrm{p},2},\ldots,s_{\mathrm{p},K}]^{\sf T}$, which is defined as $\mathbf{P}_{\mathrm{p}}=[\mathbf{p}_{\mathrm{p},1},\mathbf{p}_{\mathrm{p},2},\ldots,\mathbf{p}_{\mathrm{p},K}]\in\mathbb{C}^{N\times K},$ where $\mathbf{p}_{\mathrm{p},k}$ is the precoder for the private message of user $k$. Likewise, $\mathbf{P}_{\mathrm{r}}$ represents the precoding matrix for the vector of radar sequence symbols $\mathbf{s}_{\mathrm{r}}=[s_{\mathrm{r},1},s_{\mathrm{r},2},\ldots,s_{\mathrm{r},M}]^{\sf T}$, that is, $\mathbf{P}_{\mathrm{r}}=[\mathbf{p}_{\mathrm{r},1},\mathbf{p}_{\mathrm{r},2},\ldots,\mathbf{p}_{\mathrm{r},M}]\in\mathbb{C}^{N\times M}$, where $\mathbf{p}_{\mathrm{p},m}$ is the radar beamformer for the $m$-th radar symbol. For notational simplicity, we integrate the common and private part with a single expression, which is given by $\mathbf{P}_{\mathrm{c}}=[\mathbf{p}_{\mathrm{c}},\mathbf{P}_{\mathrm{p}}]$ and $\mathbf{s}_{\mathrm{c}}=[s_{\mathrm{c}},\mathbf{s}_{\mathrm{p}}^{\sf T}]^{\sf T}$. Lastly, we define the entire precoding vector and the symbol vector as $\mathbf{P}=[\mathbf{P}_{\mathrm{c}},\mathbf{P}_{\mathrm{r}}]$ and $\mathbf{s}=[\mathbf{s}_{\mathrm{c}}^{\sf T},\mathbf{s}_{\mathrm{r}}^{\sf T}]^{\sf T}$, respectively. The communication symbols $\mathbf{s}_{\mathrm{c}}$ and the radar sequences $\mathbf{s}_{\mathrm{r}}$ are statistically independent, as established in our previous work [2]. That is, $\mathbb{E}[\mathbf{s}_{\mathrm{c}}\mathbf{s}_{\mathrm{r}}^{\sf H}]=\mathbf{0}_{(K+1)\times M}$. In addition, the power constraint with transmit power $P$ is represented by $\mathbb{E}[{\mathbf{s}_{\mathrm{c}}\mathbf{s}_{\mathrm{c}}^{\sf H}}]=P\cdot\mathbf{I}_{K+1}$ and $\mathbb{E}[{\mathbf{s}_{\mathrm{r}}\mathbf{s}_{\mathrm{r}}^{\sf H}}]=P\cdot\mathbf{I}_{M}$, respectively. In this regard, we presume that $\|\mathbf{P}_{\mathrm{c}}\|_{\sf F}+\|\mathbf{P}_{\mathrm{r}}\|_{\sf F}\leq 1$, without loss of generality.